Question 1: Construct a one way data table to calculate the portfolio unrealized gain/loss ($) while allowing the current quote of MSFT vary from 500 to 800, incrementing by 20.

Question 2: Construct a two way data table to calculate the portfolio unrealized gain/loss ($) while allowing the current quote of MSFT vary from 500 to 800, incrementing by 20, and allowing the amount of MSFT held vary from 0 to 400, increment by 50.

Portfolio Summary
Total cost of accounts $418,003.50
Value of accounts $408,105.00
Unrealized gain/loss $ $(9,898.50)
Unrealized gain/loss % -2.37%
List of Investiments
Investment Name Quantity Purchase Price Per Share Total Cost Current Quote Market Value Gain/Loss ($) Gain/Loss (%)
AAPL 300 $698.73 $209,619.00 $490.64 $147,192.00 $(62,427.00) -29.78%
GOOG 250 $623.25 $155,812.50 $886.50 $221,625.00 $65,812.50 42.24%
MSFT 1200 $43.81 $52,572.00 $32.74 $39,288.00 $(13,284.00) -25.27%
Answer #1

Step 1: Create the spreadsheet caculating Unrealized gain/loss $

1 Portfolio Summary Question 1: One-way Data Table 3 Total cost of accounts $418,003.50 4 Value of accounts 5 Unrealized gain

Formula:

D11 =B11*C11

F11 =B11*E11

G11 =F11-D11

H11 =G11/D11

use fill handle to copy above formulas to the end of table (row 13)

B3 =SUM(D11:D13)

B4 =SUM(F11:F13)

B5 =B4-B3

B6 =B5/B3

Step 2: Create one-way data table

Write formula in cell J3 =B5

Create interval column in J4:J19 which show the current market price of MSFT.

In K3 write 1200, which indicates the Quantity of MSFT.

Select range J3:K19

On top ribbon of excel click What-if Analysis > Data Table > Row Input Cell: $B$13 ,  Column Input Cell: $E$13 > OK

One-way table is created.

($9,898.50) 1200
500 $550,813.50
520 $574,813.50
540 $598,813.50
560 $622,813.50
580 $646,813.50
600 $670,813.50
620 $694,813.50
640 $718,813.50
660 $742,813.50
680 $766,813.50
700 $790,813.50
720 $814,813.50
740 $838,813.50
760 $862,813.50
780 $886,813.50
800 $910,813.50

Question 2: Two-way data table

Write formula in cell J3 =B5

Create interval column in J4:J19 which show the current market price of MSFT.

Create interval row in K3:S3, which indicates the Quantity of MSFT.

Select range J3:S19

On top ribbon of excel click What-if Analysis > Data Table > Row Input Cell: $B$13 ,  Column Input Cell: $E$13 > OK

Two-way Data Table is created.

RI Question 2: Two-way Data Table 9 ($9,898.50 100 300 400 500$3,385.50 $26,195.00 $49,004.50$71,814.00 $94,623.50 $117,433.0